order these from least to grestest 2/5 3/4 2/3
Found 2 solutions by nerdybill, checkley77:
Answer by nerdybill(7384)

order these from least to greatest 2/5 3/4 2/3
.
First, find a common denominator -- say, 60.
Then, convert each fraction:
2/5 * 12/12 = 24/60
3/4 * 15/15 = 45/60
2/3 * 20/20 = 40/60
.
Now, comparing the "converted" fractions we see that the correct order is:
2/5, 2/3, 3/4

Answer by checkley77(12844)

2/5=.4
3/4=.75
2/3=.6667
2/5, 2/3, 3/4
.4, .667, .75